OpenStudy (hockeychick23):

You choose to participate in a raffle. Raffle tickets sell for $5 each and there are a total of 10,000 tickets. The prizes in this raffle are: 1st prize: $20,000 college scholarship (1 winner) 2nd prize: $10,000 college scholarship (2 winners) 3rd prize: $1,000 college scholarship (5 winners) 4th prize: $100 college scholarship (50 winners) You purchase one $5 raffle ticket. How much of your college tuition can you expect to be paid for by your raffle tickets purchase?

OpenStudy (hockeychick23):

@ganeshie8

OpenStudy (adi3):

ribhu do you have any idea

OpenStudy (hockeychick23):

@ganeshie8 when i calculated it i did this: (-5* 9942/10,000) + (20,000 * 1/10,000) + (10,000* 2/10,000) + (1,000*5/10,000) + (100*50/10,000)= -4.971 + 2 + 2 + .5 + .5 And this was my answer but i know its wrong so i don't no what I'm supposed to do: .029

ganeshie8 (ganeshie8):

0.029 looks good to me how do you know it is wrong ?

ganeshie8 (ganeshie8):

Ahh no wait, you forgot to subtract the ticket cost..

OpenStudy (hockeychick23):

ohhh i thought the ticket cost was just multiplied by the first answer so should i do 5-.029?

ganeshie8 (ganeshie8):

No, the random variable here is "net profit" so you need to subtract "5" from each of the prize value

OpenStudy (hockeychick23):

oh ok so i do-4.971-5 + 2-5 + 2-5 + .5-5 + .5-5?

ganeshie8 (ganeshie8):

Nope

OpenStudy (adi3):

then wht is it. I cant get it

ganeshie8 (ganeshie8):

1st prize: $20,000 college scholarship (1 winner) `(20000-5)*1/10000` 2nd prize: $10,000 college scholarship (2 winners) `(10000-5)*2/10000` 3rd prize: $1,000 college scholarship (5 winners) `(1000-5)*5/10000` 4th prize: $100 college scholarship (50 winners) `(100-5)*50/10000`

ganeshie8 (ganeshie8):

add them up : `(20000-5)*1/10000` + `(10000-5)*2/10000` + `(1000-5)*5/10000` + `(100-5)*50/10000`

OpenStudy (hockeychick23):

ok i got 1.9995+ 1.999 + .4975+.475= 4.971

ganeshie8 (ganeshie8):

also add probability of -5 : 9942*(-5/10000)

OpenStudy (hockeychick23):

ok for that i got -4.971

ganeshie8 (ganeshie8):

` (20000-5)*1/10000` + `(10000-5)*2/10000` + `(1000-5)*5/10000` + `(100-5)*50/10000` + `9942*(-5/10000)`

ganeshie8 (ganeshie8):

adding them all gives you 0 so the expected value is 0

OpenStudy (hockeychick23):

really? thats not one of my answer choices which is why I'm getting so confused, the answers are: A. -$10.20 B. $5.00 C. $4.56 D. -$3.72 E. $100

OpenStudy (hockeychick23):

@jim_thompson5910 @perl could you help me with this problem please?

OpenStudy (thomas5267):

The expected value is 0 after the inclusion of the cost of the ticket. I think your teacher made a mistake. Select B and tell him/her.